OpenStudy (anonymous):

A student’s scale recorded his weight as 160 lb. His weight is actually 164 lb. What is the percent error of the scale? A. 0.04% B. 2.4% C. 2.5% D. 0.4%

OpenStudy (anonymous):

4/160 * 100 = 1/40 * 100 = 100/40 = 10/4 = 2.50%